Two gardens. A fence of length 100 ft is to be used to enclose two


gardens. One garden is to have a circular shape, and the other to be


square.


Determine how the fence should be cut so that the sum of the areas


inside both gardens is as large as possible.

1
Expert's answer
2021-12-07T06:39:34-0500

Let side of the square garden is a and radius of the circular garden is r.

We have to maximize "a^{2}+\pi r^{2}"

 "\begin{aligned}

&\text { Subject to, } 4 a+2 \pi r=100 \\

&\text { i.e., } 2 a+\pi r=50

\end{aligned}"

The Lagrange function "L=\left(a^{2}+\pi r^{2}\right)+\lambda(2 a+\pi r-50)."

"\therefore \begin{aligned}

\frac{\partial L}{\partial a} & \equiv 2 a+2 \lambda=0 \quad \Rightarrow \quad \lambda=-a \\

\frac{\partial L}{\partial r} & \equiv 2 \pi r+\lambda \pi=0 \quad \Rightarrow \lambda=-2 r .

\end{aligned}"

"\begin{aligned}

\therefore & 2(-\lambda)+\pi\left(-\frac{\lambda}{2}\right)=50 \\

& \Rightarrow \lambda\left(-2-\frac{\pi}{2}\right)=50 \\

& \Rightarrow \lambda=\frac{-100}{4+\pi} \\

\therefore \quad & a=\frac{100}{4+\pi} \\

\text { and, } & r=\frac{50}{4+\pi}

\end{aligned}"

"\therefore" The required fence for square "=\frac{400}{4+\pi} \mathrm{ft}" and for circular garden "=\frac{100 \pi}{4+\pi} f_{t}."